# Stepwise Reasoning Error Disruption Attack of LLMs

### v1

@@ 3.3 Parameter Analysis
In the SEED attack, $\sigma$ is the hyperparameter that controls the proportion of injected reasoning steps, which intuitively influences the attack performance. To explore its impact, we evaluate the performance under different values of $\sigma$. The results are shown in Figure 4. We observe that the performance varies across different models and tasks. In most cases, a range between 0.4 and 0.6 yields competitive performance. When $\sigma$ is lower, fewer reasoning steps are injected. Consequently, the target LLM pays less attention to the injected steps and is more inclined to rely on its original reasoning process, resulting in a significant drop in ASR.
On the other hand, when $\sigma$ is higher, the ASR also drops noticeably in some cases, particularly with GPT4-o and Qwen-2.5 on MATH. We hypothesize that the reason behind this behavior is that over-injecting reasoning steps sometimes makes the LLM more robust. When too many incorrect steps are injected, the model may become aware of the inconsistencies and try to correct them, leading to a more cautious approach. This increased awareness might cause the LLM to rely more on its internal reasoning, which could prevent the attack from being as effective. More results are provided in Appendix C due to space limitation.
3.4 Ablation Study
Two key components of SEED-S are the prepending of a wrong answer and the 2-stage reasoning step generation, which involves: 1) solving the raw problem to generate the correct solution, and 2) in multiple-choice tasks, selecting a different answer and generating a corresponding solution with reasoning steps that lead to the selected answer. For open-ended tasks, the solution is directly created with reasoning steps that lead to the incorrect answer, without the need to choose a different answer.
Figure 5 illustrates the impact of these components, showing that both contribute to the overall performance. Notably, on CSQA, the 2-stage generation has a more significant effect, as in multiple-choice tasks, the LLM tends to notice when the final answer is not among the provided answer choices, prompting it to correct the error. The 2-stage reasoning generation ensures alignment between the given answer choice and the generated solution, specifically in multiple-choice tasks. More results are presented in Appendix C.
Table 4: ASR of SEED-P on questions answered correctly and incorrectly without the attack. Raw_C represents the attack performance on correctly answered questions, while Raw_I denotes the performance on incorrectly answered questions. Highest ASR are highlighted within each model for a given dataset setting.
@@ 4 Related Work
4.1 Prompt-based Reasoning of LLMs
Enhancing the reasoning capabilities of large language models (LLMs) remains a prominent research focus Yang et al. (2024); Ning et al. . Among various approaches, the Chain of Thought (CoT) paradigm has proven particularly effective, as shown by foundational works like Wei et al. (2022) and Kojima et al. (2022). These studies demonstrated that integrating explicit reasoning steps—via reasoning exemplars or step-by-step instructions—significantly boosts LLM reasoning performance. Subsequent research has refined CoT with methods like the self-consistency strategy Wang et al. , which employs majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage approach that decomposes problems before solving sub-tasks. Moreover, techniques like Automatic and semi-automatic Prompting enhance CoT by optimizing prompts Zhang et al. ; KaShun et al. ; Pitis et al. (2023). Extensions to more complex structures, such as trees Yao et al. (2024) and graphs Besta et al. (2024), further expand CoT’s potential.
@@ 4.2 Prompt-based Attack on LLMs
A key area of research aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024a); Zhao et al. (2023), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature.
Figure 5: Ablation study of SEED-P on CSQA dataset. w/o WA: without wrong answer, w/o 2-stage: without 2-stage reasoning generation. Results show both components are important, especially 2-stage generation.
@@ 5 Conclusion and Future Works
We propose Stepwise Reasoning Error Disruption attack (SEED), a novel method targeting LLMs’ reasoning capabilities by injecting misleading steps with deliberate errors to disrupt their reasoning process. Through experiments on four datasets and LLMs, we demonstrate our method’s effectiveness with two variations, achieving high success rates while remaining stealthy. Our attack reveals LLMs’ vulnerability to adversarial reasoning steps, especially in multi-turn scenarios where early errors can cascade through the reasoning chain. Our findings highlight the need for robust defenses to protect LLMs’ reasoning integrity in critical applications.
@@ 6 Limitation
Additionally, our attack method may inadvertently generate potentially harmful or offensive content in the output solutions for the modified questions. This risk arises due to the nature of adversarial attack, which alter the model’s responses in unintended ways. Without rigorous safeguards, including targeted controls and thorough examination of outputs, the potential for generating inappropriate or harmful content cannot be fully mitigated. Future efforts should focus on integrating content moderation techniques and ethical safeguards to minimize these risks while maintaining the effectiveness of the attack method.
Appendix A Details for the Metric
Attack Successful Rate. The Attack Success Rate (ASR) measures the proportion of originally correct answers that become incorrect after the attack is applied:
Here, $C_{\text{original}}$ represents the set of correctly answered questions before the attack, and $W_{\text{attack}}$denotes the set of wrongly answered questions after the attack.This metric serves as a direct and quantitative indicator of the effectiveness of the attack in compromising the model’s reasoning capabilities.
Appendix B Details for the Datasets
@@
Figure 6: Prompt utilized for SEED-S and SEED-P attack, and attack detection.
Appendix C More Experiment Results
Due to space constraints, additional results from the parameter analysis and ablation study are presented in Figure 10 and Figure 7.
@@ Appendix C More Experiment Results
Appendix D Implementation of SEED attack
@@
Appendix E Evaluation of SEED Attack Transferability
We evaluate the transferability of the SEED attack across different datasets by conducting attacks using various LLMs on a target LLM, with the results shown in Figure 8. The results reveal that the proposed SEED attack consistently achieves a high ASR across diverse assistant and target LLM combinations, highlighting its stability and effectiveness. Furthermore, Qwen and GPT-4 stand out as the most robust target LLMs, showing relatively strong resistance to attacks from different sources. On the other hand, GPT-4 exhibits the most potent attacking capability, outperforming other models against nearly all target LLMs across datasets, especially on the CSQA dataset. This dual strength underscores GPT-4’s exceptional performance in both offensive and defensive roles.
@@ Appendix E Evaluation of SEED Attack Transferability
a) Performance on MATH dataset
b) Performance on CSQA dataset

### v3

@@ 3.3 Parameter Analysis
In the SEED attack, $\sigma$ is the hyperparameter that controls the proportion of injected reasoning steps, which intuitively influences the attack performance. To explore its impact, we evaluated the performance of SEED-P under different values of $\sigma$. The results, shown in Figure 3, indicate that performance varies across different models and tasks. Generally, a $\sigma$ range between 0.4 and 0.6 yields competitive performance. Lower $\sigma$ values result in fewer injected reasoning steps, causing the target LLM to rely more on its original reasoning process and leading to a significant drop in ASR.
Conversely, higher $\sigma$ values also cause noticeable ASR drops in some cases, particularly with GPT-4o and Qwen-2.5 on MATH. We hypothesize that over-injecting reasoning steps can make the LLM more robust. When too many prior steps are introduced, the LLM focuses more on reviewing its prior reasoning rather than continuing with subsequent inference. This increased scrutiny helps the LLM detect inconsistencies and attempt corrections, leading to a more cautious reasoning approach and reducing the attack’s effectiveness. Additional results are provided in Appendix I due to space limitations.
Table 4: MSR of SEED-P on questions answered correctly and incorrectly without the attack. Raw_C represents the attack performance on correctly answered questions, while Raw_I denotes the performance on incorrectly answered questions. Highest MSR are highlighted within each model for a given dataset setting.
@@ 4 Related Work
4.1 Reasoning of LLMs
Enhancing reasoning in large language models (LLMs) remains a key research focus Yang et al. (2024); Ning et al. . The Chain of Thought (CoT) paradigm has been particularly effective, as shown by Wei et al. (2022) and Kojima et al. (2022), demonstrating that explicit reasoning steps, such as exemplars or step-by-step instructions, improve LLM performance. Subsequent work refined CoT with techniques like self-consistency Wang et al. , which uses majority voting across reasoning paths, and Least-to-Most Zhou et al. , a two-stage problem decomposition approach. Further extensions to trees Yao et al. (2024) and graphs Besta et al. (2024) expand CoT’s capabilities. Recent advances in long reasoning methods require LLMs to iteratively build upon prior steps, facilitating reflection Madaan et al. (2024); Zhao et al. (2024) or tree search Guan et al. (2025); Zhang et al. (2024) for subsequent reasoning steps, further expand the reasoning ability of LLMs. This reliance on step-by-step reasoning, however, raises new concerns regarding the vulnerability of LLMs.
@@ 4.2 Prompt-based Attack on LLMs
A key area of research aimed at ensuring the safety and robustness of LLMs involves developing methods to attack these models, prompting the generation of undesirable content Deng et al. (2023); Chu et al. (2024); Yu et al. (2024). One prominent category within this field focuses on "jailbreak" attacks, which bypass alignment mechanisms to elicit harmful or unsafe outputs Yi et al. (2024); Mehrotra et al. (2023); Zheng et al. (2024). However, our work is not directly related to jailbreak attacks. Instead, we focus on adversarial attacks, which subtly manipulate outputs without noticeable input modifications Xu et al. (2022); Kandpal et al. ; Xu et al. . While earlier studies targeted traditional NLP tasks such as sentiment analysis and classification Wang et al. (2024); Zhao et al. (2023), recent efforts have increasingly focused on attacking LLM reasoning processes Xiang et al. (2024); Xu et al. (2024). BadChain leverages backdoor vulnerabilities by embedding triggers within in-context learning demonstrations, but its applicability remains limited to specific contexts Xiang et al. (2024). Moreover, a critical drawback of BadChain is its nearly 100% detection rate, rendering it unsuitable for practical deployment. Similarly, UPA and MPA methods proposed by Xu et al. (2024), which instruct LLMs to generate answers before reasoning, often yield outputs that are easily identifiable, compromising their covert nature. Therefore, these approaches struggle to strike an effective balance between attack potency and stealth.
@@ 5 Conclusion and Future Works
We propose Stepwise Reasoning Error Disruption attack (SEED), a novel method targeting LLMs’ reasoning capabilities by injecting misleading steps with deliberate errors to disrupt their reasoning process. Through experiments on four datasets and LLMs, we demonstrate our method’s effectiveness with two variations, achieving high success rates while remaining stealthy. Our attack reveals LLMs’ vulnerability to adversarial reasoning steps, especially in multi-step reasoning scenarios where early errors can cascade through the reasoning chain. Our findings highlight the need for more robust defenses to protect LLMs’ reasoning integrity.
@@ 6 Limitation
Additionally, our attack method may inadvertently generate potentially harmful or offensive content in the output solutions for the modified questions. This risk arises due to the nature of adversarial attack, which alter the model’s responses in unintended ways. Without rigorous safeguards, including targeted controls and thorough examination of outputs, the potential for generating inappropriate or harmful content cannot be fully mitigated. Future efforts should focus on integrating more complicated content moderation techniques and ethical safeguards to minimize these risks while maintaining the effectiveness of the attack method.
Appendix A Details for the Datasets
@@
Figure 6: The form used for human evaluation of coverage of each attack approach.
Appendix B Implementation of SEED attack
@@
Appendix C Details for the Metric
Attack Success Rate. The Attack Success Rate (ASR) measures the proportion of originally correct answers that become incorrect after the attack is applied:
where $C_{\text{original}}$ represents the set of correctly answered questions before the attack, and $W_{\text{attack}}$ denotes the set of wrongly answered questions after the attack. This metric serves as a direct and quantitative indicator of the effectiveness of the attack in compromising the model’s reasoning capabilities.
Modification Successful Rate The Modification Successful (MSR) quantifies the proportion of problems that are altered by the attack:
$MSR=\frac{|W_{\text{attack}}|}{|C_{\text{original}}\cup I_{\text{original}}|},$
where $I_{\text{original}}$ represents the set of incorrectly answered questions before the attack.
Appendix D Human Evaluation on Covert Detection
We conducted a comprehensive human evaluation study with 20 PhD students in computer science, using the MATH and CSQA datasets. Each evaluator independently assessed 10 solutions for each attack scenario, along with 10 unattacked solutions, resulting in a total of 60 evaluations per dataset. To ensure a focused and consistent evaluation, we instructed each evaluator to spend no more than 10 seconds on each solution, emphasizing the coherence of the Chain of Thought (CoT) rather than the correctness of the reasoning and calculations. The evaluation form used is shown in Figure 6.
Table 5: Human evaluation on covert detection.
BadChain MPA UPA SEED-S SEED-P Pure
MATH 0.97 0.44 0.36 0.17 0.20 0.09
CSQA 0.96 0.42 0.38 0.15 0.21 0.08
The evaluators’ assessments closely aligned with the results from our GPT-4o model, as detailed in Table 5. BadChain achieved a perfect 100% detection rate, while UPA and MPA demonstrated detection rates of 37% and 43%, respectively. In contrast, SEED-S and SEED-P showed relatively low detection rates of 16% and 21%, which were close to the 9% baseline observed for unattacked solutions.
Each PhD student spent approximately half an hour on the evaluation process, for which they were compensated US\$15 per evaluation, yielding an hourly wage of at least US\$30.
Figure 7: Case study on SEED-S/P attack. The red font highlights the modified content.
Appendix E Case Study
As shown in Figure 7, in SEED-S, $LLM_{assist}$ automatically makes modifications based on different types of problems. For instance, in mathematical problems, it modifies the intermediate calculation steps, while in multiple-choice reasoning tasks, it analyzes the options with varying degrees of inclination. However, since it can only modify one step at a time, it may not always be sufficient to persuade the LLM to output the target result. In SEED-P, $LLM_{assist}$ typically adjusts numerical values in math problems, while for common-sense reasoning tasks, it automatically identifies and modifies the most influential elements, often verbs or nouns, that affect the final outcome.
Appendix F Evaluation of SEED Attack Transferability
We evaluate the transferability of the SEED attack across different datasets by conducting attacks using various LLMs on a target LLM, with the results shown in Figure 9. The results reveal that the proposed SEED attack consistently achieves a high ASR across diverse assistant and target LLM combinations, highlighting its stability and effectiveness. Furthermore, Qwen and GPT-4o stand out as the most robust target LLMs, showing relatively strong resistance to attacks from different sources. On the other hand, GPT-4o exhibits the most potent attacking capability, outperforming other models against nearly all target LLMs across datasets, especially on the CSQA dataset. This dual strength underscores GPT-4o’s exceptional performance in both offensive and defensive roles.
Table 6: Detailed results of prompt-based self-review mitigation against SEED-P attack under zero-shot setting.
MATH GSM8K CSQA MATHQA
Llama3 SEED-P 0.514 0.425 0.626 0.518
Mitigation 0.508 0.418 0.625 0.508
Qwen SEED-P 0.447 0.418 0.384 0.346
Mitigation 0.44 0.406 0.378 0.344
Mistral SEED-P 0.722 0.804 0.767 0.759
Mitigation 0.685 0.724 0.698 0.744
GPT-4o SEED-P 0.286 0.191 0.605 0.45
Mitigation 0.276 0.184 0.568 0.432
Appendix G Prompt-based mitigation
We thoroughly tested prompt-based self-review mitigation under zero-shot setting by appending “review your reasoning steps before providing final answer” to the prompt. Our detailed results shown in Table 6 reveals modest improvement, suggesting that straightforward prompt-based defenses may require enhancement to effectively counter SEED-P attack.
Appendix H Ablation Study
Two key components of SEED-P are the prepending of a wrong answer and the 2-stage reasoning step generation, which involves: 1) solving the raw problem to generate the correct solution, and 2) in multiple-choice tasks, selecting a different answer and generating a corresponding solution with reasoning steps that lead to the selected answer. For open-ended tasks, the solution is directly created with reasoning steps that lead to the incorrect answer, without the need to choose a different answer. In the absence of the two-stage process, the LLM directly modifies the question rather than first generating the correct answer and subsequently selecting an incorrect answer for reasoning.
Figure 5 illustrates the impact of these components, showing that both contribute to the overall performance. Notably, on CSQA, the 2-stage generation has a more significant effect, as in multiple-choice tasks, the LLM tends to notice when the final answer is not among the provided answer choices, prompting it to correct the error. The 2-stage reasoning generation ensures alignment between the given answer choice and the generated solution, specifically in multiple-choice tasks.
Appendix I More Experiment Results
Due to space constraints, additional results from the parameter analysis are presented in Figure 8 .
a) Performance on MATH dataset
b) Performance on CSQA dataset